if tangent to the curve root of x+root of y=root of a at a point on it cuts the axes ox and oy at p ang q respectively then op+oq is
Options~
A 2a B a
C 1/2a D none of these

Grade:12

2 Answers

Bevkoof Singh
10 years ago

since

root x+root y=root a

so coordinates when this curve cut x axis

(a,0)

simil. for y

(0,a)

so p=a=q

thus ans. is (A) 2a

Raj
4 years ago
√y= √a -√x
Now diff. w.r.t to x
dy/dx= - √y/√x
y - y'/x - x' = -√y/√x
Now if it forms triangle then PO^2+OQ^2= PQ^2
Therefore OP =OQ = a 
PQ = 2a